# OneRec-V2 Technical Report

### v1

@@ Abstract
1. Lazy Decoder-Only Architecture: A streamlined, decoder-only design that eliminates encoder bottlenecks and simplifies cross-attention, reducing total computation by 94% and training resources by 90% (see Figure 1 right). This efficiency enables the successful scaling of the model to 8B parameters, with a consistent loss decrease observed throughout the process (see Figure 1 left).
@@ 2.3.4 Model Scaling
The convergence loss demonstrates consistent improvements with increased model size, decreasing from 3.57 for the 0.1B model to 3.19 for the 8B model. The most substantial improvements occur in the sub-billion parameter regime, with a loss reduction of 0.3 when scaling from 0.1B to 1B parameters. The marginal gains diminish, with the 4B model achieving a further reduction of only 0.03 compared to the 2B model, indicating that scaling beyond 2B parameters remains challenging.
@@ 3.1.2 Reinforcement Learning
However, these methods do not provide a complete and comprehensive consideration of gradient stability. In particular, for negative samples, the absence of an upper bound on the policy ratio can easily lead to gradient explosion, causing the model’s performance to collapse. Dual-clip applies an upper bound truncation to the policy ratio for negative samples. While this improves stability, it discards too many negative samples, which slows the convergence. In OneRec-V1 (Zhou et al., 2025), we propose Early Clipped GRPO (ECPO), which truncates the gradient upper bound for negative samples, thereby retaining more samples while enhancing stability.
@@ 5 Conclusion, Limitations, and Future Directions
    Scaling: Although we observed a continuous decrease in loss as the model scaled from 0.1B to 8B, the downward trend does not strictly adhere to scaling laws (Kaplan et al., 2020). This indicates that model scaling requires further exploration, and we will continue to invest in research on aspects such as data organizations, model architectures, and pre-training methods.

### v4

@@ Abstract
1. Lazy Decoder-Only Architecture: A streamlined, decoder-only design that eliminates encoder bottlenecks and simplifies cross-attention, reducing total computation by 94% and training resources by 90% (see Figure 1 right). This efficiency enables the successful scaling of the model to 8B parameters and, notably, the convergence loss closely follows the empirical scaling law. As the model scales, we observe a smooth and predictable decrease in loss consistent with the scaling law fit (see Figure 1 left, and Figure 6).
@@ 2.3.4 Model Scaling
Figure 6: Dense Model Scaling Law Curve. The scaling law is fitted by $\hat{L}(N)\triangleq E+\frac{A}{N^{\alpha}}$, where $E=3.13$, $A=3660$, $\alpha=0.489$.
@@
Our empirical results show reasonable agreement with the theoretical scaling law. While the general Chinchilla scaling law expression is
$\hat{L}(N,D)\triangleq E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$
where $N$ is the number of model parameters and $D$ is the number of training tokens Hoffmann et al. (2022), our experimental setup keeps $D$ fixed. In this case, the term $\frac{B}{D^{\beta}}$ is a constant and can be absorbed into the intercept $E$. Therefore, for fixed data, the scaling law simplifies to
$\hat{L}(N)\triangleq E^{\prime}+\frac{A}{N^{\alpha}}$
where $E^{\prime}=E+\frac{B}{D^{\beta}}$. Our empirical results fit this functional form very well: $E=3.13$, $A=3660$, $\alpha=0.489$, as shown in Figure 6.
@@ 3.1.2 Reinforcement Learning
However, these methods do not provide a complete and comprehensive consideration of gradient stability. In particular, for negative samples, the absence of an upper bound on the policy ratio can easily lead to gradient explosion, causing the model’s performance to collapse. Dual-clip (Ye et al., 2020) applies an upper bound truncation to the policy ratio for negative samples. While this improves stability, it discards too many negative samples, which slows the convergence. ECPO (Zhou et al., 2025) mitigates gradient explosion on negative samples by applying clipping directly to $\pi_{old}$ rather than to the ratio $\pi_{\theta}/\pi_{old}$. This strategy retains a larger proportion of training samples while improving optimization stability. Similarly, CISPO (Chen et al., 2025) and GPPO (Su et al., 2025) adopt related techniques to keep the ratio within a reasonable range, while preserving gradient signals from more samples. In OneRec V1, we employ ECPO (Early Clipped GRPO), formally defined as:
@@ 4 Online A/B Test
To further validate our findings, we conducted an additional experiment with caching disabled where all traffic in a separate 1% experimental group requests OneRec-V2 (detailed results in Appendix D). This comprehensive evaluation confirms the substantial improvements in user engagement metrics, with interaction indicators such as likes, follows, comments, and forwards showing remarkable gains of 9.6% to 29.2% across platforms. While these results demonstrate OneRec-V2’s strong performance in driving user engagement, they also reveal important ecosystem-level considerations, including significant reductions in cold-start video views (44.7% and 36.7% for Kuaishou and Kuaishou Lite respectively) and increased cluster density.
@@ 5 Conclusion, Limitations, and Future Directions
    Scaling: As the model scaled from 0.1B to 8B parameters, we observed a consistent decrease in loss, which is remarkably well described by the empirical scaling law proposed by Hoffmann et al. (2022) (Hoffmann et al., 2022). Our results show excellent agreement with this scaling relation, as evidenced in Figure 6. This validates the effectiveness of the chosen architecture and indicates that further scaling and architectural innovations can potentially yield even better performance.